Amortization Schedule


Year


Cash


Interest

Amount
Unamortized

Carrying
Value

1/1/2014

$54,103 $ 161,497

2014

$21,560 $24,225 51,438 164,162

2015

21,560 24,624 48,374 167,226

2016

21,560 25,084 44,850 170,750

2017

21,560 25,613 40,797 174,803

2018

21,560 26,220 36,137 179,463

2019

21,560 26,919 30,778 184,822

2020

21,560 27,723 24,615 190,985

2021

21,560 28,648 17,527 198,073

2022

21,560 29,711 9,376 206,224

2023

21,560 30,936 215,600

(c) Determine the stated interest rate and the effective-interest rate. (Round answers to 0 decimal places, e.g. 18%.)
(d) On the basis of the schedule above, prepare the journal entry to record the issuance of the bonds on January 1, 2014. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)
(e) On the basis of the schedule above, prepare the journal entry to reflect the bond transactions and accruals for 2014. (Interest is paid January 1.) (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)
(f) On the basis of the schedule above, prepare the journal entries to reflect the bond transactions and accruals for 2021. Novak Corporation does not use reversing entries. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)
Answer #1

c) Stated rate of interest = 21560/215600 = 10%

Market rate of interest = 24225/161497 = 15%

d) Journal entry

Date General Journal Debit Credit
Jan 1 Cash 161497
Discount on bonds payable 54103
Bonds payable 215600

e) Journal entry

Date General Journal Debit Credit
Dec 31 Interest expense 24225
Discount on bonds payable 2665
Interest payable 21560

f) Journal entry

Date General Journal Debit Credit
Jan 1,2021 Interest payable 21560
Cash 21560
Dec 31, 2021 Interest expense 28648
Discount on bonds payable 7088
Interest payable 21560